# LLM Social Simulations Are a Promising Research Method

### v1

@@ 4.5.2 Iterative evaluation
There is much that simulation-focused researchers can do for each of these five challenges, but ultimately, research efforts in other domains will likely reshape the possibility landscape, as seen in the stunning advances in LLM capabilities in recent years. Therefore, our vision for sims centers on iterative evaluation so that the nascent sims community can keep pace.
While there is divergence between human and LLM behavior, as discussed in Section 3.4, there are reasons to expect AI to become more humanlike over time, mitigating the alienness challenge. Some have argued that AI world-models will converge to one another (Huh et al., 2024). Digital minds may also converge in behavior and mechanisms to human minds as both are optimized towards accurate and efficient models of reality, whether through evolutionary pressure or an artificial learning paradigm. Behavioral convergence has been observed in many areas, such as LLMs exhibiting humanlike value trade-offs (Liu et al., 2024c) and humanlike failures of “overthinking” when allowed to “reason” through token generation before arriving at a final answer (Liu et al., 2024b). While mechanistic divergence is more difficult to identify, Binz et al. (2024) found that their LLM fine-tuned on social science data, Centaur, has internal representations that more consistently predict human whole-brain fMRI data than those of the model on which it is based, Llama-3.1-70B. Both Centaur and Llama-3.1-70B were found to predict fMRI data better than a traditional cognitive model.
@@
Likewise, there are reasons to expect that this reduction in alienness as well as scaling and algorithmic advances will begin to address the challenge of generalization. In particular, LLMs and other AI systems are increasingly able to utilize vast datasets, such as large-scale purchasing data from the world’s largest companies to understand economic behavior (Einav & Levin, 2014) or location tracking data from millions of mobile devices to understand human movement and travel (Li et al., 2024). Even if one views LLMs as exclusively interpolating between existing data points, existing data sources could grow to encompass much of the human behavior social scientists aim to explain.
@@
The most straightforward method to evaluate OOD performance is to use data that is, from the perspective of the model, not yet incorporated in human knowledge. Three studies in our review tested sims on predictions of novel data that was past the model’s training cutoff, data that was not yet collected until after the simulation, and data in holdout sets that was not yet publicly accessible, as detailed in Section A.2.4. It will be important to compare the predictions of LLMs on novel data with human predictions of the same data. Of the studies we reviewed, Hewitt et al. (2024) was the only study that did so. They compared LLM accuracy with the accuracy of laypeople (crowdworkers on Prolific) and experts (data collected by authors of some of the original studies). On their adjusted correlation measure, laypeople predicted a large majority of variation (84%), GPT-4 outperformed the laypeople (91%), GPT-3.5 slightly underperformed (82%), and two GPT-3 versions greatly underperformed (-9% and 25%). Expert forecasts were only available for some studies, but they tended to perform about as well as GPT-4. However, Hewitt et al. did not report how well the sims performed specifically on the subset of data that humans found difficult to predict, which is the subset that would be most fruitful for scientific discovery.
@@
We have not yet seen, but we encourage, preregistration of LLM simulation predictions, particularly if preregistration includes human expert forecasts. That way, researchers can directly observe how sims compare to novel data. If those preregistrations are shared publicly, the research community can mitigate publication bias. Once this validation is successful in a particular domain, such as the type of survey experiments studied by Hewitt et al. (2024), researchers should feel more comfortable using LLM social simulations, at least for exploratory research. In any case, it will be important to map out the areas in which LLM simulation is more or less successful at generalization. The most challenging areas may be related to each other in ways amenable to particular methods.
@@ 5 Alternative views
We have argued that LLM social simulations currently have substantial issues but also that there are promising future directions. Because this is a new research area, there has been limited discussion of such claims, but many of the reviewed works emphasize significant challenges and the limits of initial results (e.g., Gao et al., 2024; Wang et al., 2024a, b). For example, in their commentary, Agnew et al. (2024) reviewed some of the recent LLM simulation proposals, concluding that they “ignore and ultimately conflict with foundational values of work with human participants” due to problems such as diversity.
More generally, LLMs have been called “stochastic parrots” (Bender et al., 2021) that are “fundamentally not like us” (Shanahan, 2022) with “ineradicable defects” (Chomsky et al., 2023), among other critiques (Bender & Koller, 2020; LeCun, 2023; Marcus, 2022). These views imply that LLM capabilities are deeply limited, particularly on social simulation—a task that inherently requires the system to be humanlike or to understand humanlikeness sufficiently to make accurate predictions. We believe that theoretical arguments on this topic can only tell us so much, and we provide more detail on the limitations of our work in Appendix C. Ultimately, we believe that it will only be possible to validate the promise of LLM social simulations through rigorous empirical testing.
@@ 6 Conclusion
By developing an agenda for LLM social simulations or “sims” based on five tractable challenges, we show that the difficulties are not as insurmountable as some researchers have suggested. There is inevitably much uncertainty in the future of AI progress, but the rapid pace of technical advances requires us to take a step back and consider whether this endeavor could succeed. We argue that, by thoughtfully incorporating contemporary methods and future AI capabilities, sims are a promising research method to accelerate social science, open up new areas of inquiry into human behavior, and generate humanlike synthetic data to support the development of safe and beneficial AI.
7 Acknowledgments
We are grateful for feedback and support from Suhaib Abdurahman, Sophia Kazinnik, Jiaxin Pei, Joon Sung Park, and other colleagues in the University of Chicago Knowledge Lab, Stanford Human-Computer Interaction (HCI), Stanford Natural Language Processing (NLP), the Stanford Institute for Human-Centered Artificial Intelligence (HAI), and the Stanford Digital Economy Lab (DEL).
@@ A.1.2 Fundamental limitations
There are fundamental problems that seem intractable even with an influx of funding. If researchers hope to collect real-world data, many of the most interesting research questions, such as the psychology of world leaders, the effects of large-scale policy change, or the effects of large-scale events on the general public (e.g., during a bank run; Kazinnik, 2023), would be logistically infeasible with any realistic amount of resources. Many social science questions are not even about real-world data, including future possibilities or past counterfactuals of the form “What would have happened if…?” Natural science has struggled less with these issues because the strictness of physical laws and the homogeneous nature of physical entities allow for reliable laboratory testing, such as identical copies of a molecule that are each subjected to different conditions. Human behavior has no chemical formula that can be isolated and studied in a lab.
@@ A.2 The promise of LLM simulations
This humanlikeness suggests new paradigms of where AI could be applied. Over seven decades since Alan Turing famously asked “Can machines think?” and proposed the Turing test (Turing, 1950), researchers at frontier AI labs now claim that current AI “attains a form of general intelligence” (Bubeck et al., 2023) and has reached “Level 1 General AI (‘Emerging AGI’)” (Morris et al., 2024). Generating realistic human behavioral data is the quintessential task that all humans can do, for at least one human—themselves. Thus, it would be reasonable to expect that in the next phase of AI systems, they would be able to do the same. As LLM pretraining is based on next-token prediction using trillions of words of human-written text, it is natural to expect high performance in simulating human behavior, at least that of reading and writing.
@@ A.2.3 Interview-based prompt variation
Park et al. (2024a) conducted two-hour interviews with 1,052 participants, prepending them one at a time to the study information to create prompts for GPT-4o that would simulate each participant. To evaluate this method, they conducted an extensive battery of popular social science studies with each participant twice, two weeks apart. The difference between the human participant’s responses over the two weeks was used as the ceiling for simulation accuracy. Most strikingly, they found that the simulations were 85% as accurate in predicting responses to the General Social Survey [^5] https://gss.norc.org/ as participants’ first test results were at predicting their second test results.
@@ Appendix C Limitations of Our Work
The literature on LLM simulations has only recently emerged and is scattered across disciplines and venues with terminology and framing, so we were unable to identify work on the topic through conventional literature review methods (e.g., querying Scopus or Web of Science databases). We expect that we identified a majority of published work and preprints on the topic, given the consistency with which the authors have tended to become aware of new work, but this remains a concern—particularly for work that would be bibliometrically cut off from the aforementioned papers, such as if those unknown papers have not cited these works, if they are in different scholarly venues, or if they appear in non-English languages.
It is also difficult to draw a clear line between this topic and related topics, such as using LLMs for text annotation. For example, Veselovsky et al. (2023a) explicitly focus on the distributional accuracy of LLM-generated data relative to human data. The human data were from Abu Farha et al. (2022), social media posts made by crowdworkers that they self-identified as “sarcastic.” While Veselovsky et al. (2023a) have clearly compared LLM-generated and human-generated data, the data itself is more used for natural language processing research than for social science research with human subjects, so it has clear but limited relevance to our work. Likewise, our focus has been the simulation of humans in isolation, such as a survey context, but LLMs can also be used for multi-agent and multi-turn contexts (e.g., Park et al., 2023; Piao et al., 2025). We maintain this focus in part because the success of more complex simulations will likely rely on the ability to simulate each human response in isolation.

### v2

@@ 4.5.2 Iterative Evaluation
There is much that simulation-focused researchers can do to overcome these five challenges, but ultimately, the largest changes may come from more general advances in LLM capabilities. Therefore, we see it as essential to build methods of iterative evaluation to make the best use of advancements as soon as they arrive. This includes toolkits of simulation methods, such as different approaches to fine-tuning (Binz et al., 2024; Suh et al., 2025), and datasets, such as Psych-101 (Binz et al., 2024), SubPOP (Suh et al., 2025), and Twin-2K-500 (Toubia et al., 2025).
While there is divergence between human and LLM behavior, as discussed in Section 3.4, there are reasons to expect AI to become more humanlike over time, mitigating the alienness challenge. Some have argued that AI world-models will converge to one another (Huh et al., 2024). Digital minds may also converge in behavior and mechanisms to human minds as both are optimized towards accurate and efficient models of reality, whether through evolutionary pressure or an artificial learning paradigm. Behavioral convergence has been observed in many areas, such as LLMs exhibiting humanlike value trade-offs (Liu et al., 2024c) and humanlike failures of “overthinking” (Liu et al., 2024b).
Mechanistic divergence is more difficult to identify, but Binz et al. (2024) found that their LLM fine-tuned on social science data, Centaur, has internal representations that more consistently predict human whole-brain fMRI data than those of the model on which it is based, Llama-3.1-70B. Both Centaur and Llama-3.1-70B were found to predict fMRI data better than a traditional cognitive model.
@@
Likewise, there are reasons to expect that this reduction in alienness as well as scaling and algorithmic advances will begin to address the challenge of generalization. In particular, LLMs and other AI systems are increasingly able to utilize vast datasets, such as large-scale purchasing data from the world’s largest companies to understand economic behavior (Einav & Levin, 2014) or location tracking data from millions of mobile devices to understand human movement and travel (Li et al., 2024). Even if one views LLMs as exclusively interpolating between existing data points, the space covered by interpolation could grow to encompass much of the human behavior social scientists aim to explain, even behavior not yet been captured in scientific datasets or social theory.
@@
The most straightforward method to evaluate OOD performance is to use data that is not yet incorporated in training data. Three studies in our review tested sims on predictions of novel data that was past the model’s training cutoff, data that was not yet collected until after the simulation, and data in holdout sets that was not yet publicly accessible, as detailed in Section A.2.4. It will be important to compare the predictions of LLMs on novel data with human predictions of the same data.
Of the studies we reviewed, Hewitt et al. (2024) was the only study that did so. They compared LLM accuracy with the accuracy of laypeople (crowdworkers on Prolific) and experts (data collected by authors of some of the original studies). On their adjusted correlation measure, laypeople predicted a large majority of variation (84%), GPT-4 outperformed the laypeople (91%), GPT-3.5 slightly underperformed (82%), and two GPT-3 versions greatly underperformed (-9% and 25%). Expert forecasts were only available for some studies, but they tended to perform about as well as GPT-4. However, Hewitt et al. did not report how well the sims performed specifically on the subset of data that humans found difficult to predict, which is the subset that would be most fruitful for scientific discovery.
@@
We have not yet seen, but we encourage, preregistration of LLM simulation predictions, particularly if preregistration includes human expert forecasts so that researchers can directly observe how sims compare to novel data. If those preregistrations are shared publicly and with verification they are from before data collection, then the research community can mitigate publication bias. Once this validation is successful in a particular domain, such as the type of survey experiments studied by Hewitt et al. (2024), researchers should feel more comfortable using LLM social simulations, at least for exploratory research. In any case, it will be important to map out the areas in which LLM simulation is more or less successful at generalization. The most challenging areas may be related to each other in ways amenable to particular methods.
5 Applications
Figure 1: Six applications of LLM social simulations. The most difficult applications are complete studies that are human-possible (HP), where it would be possible to use human subjects, or human-impossible (HI), such as large-scale policy experiments.
We believe that sims can already be used for pilot studies, preliminary runs to surface possible issues and estimate effect sizes for a complete study with human subjects. These results would be primarily used to inform methodological choices, rather than the underlying research topic. Current sims can be used for exploratory studies—oriented towards theory-building, brainstorming, and sketching out possibilities of future studies. Errors in exploratory studies typically have lower costs than errors in confirmatory studies.
More ambitious applications include exact replication of a human subjects study to increase or decrease confidence in results, possibly with an analysis of the combined dataset (e.g., Broska et al., 2025). Moreover, studies often require a multitude of methodological choices—some of which can lead to opposite conclusions (Breznau et al., 2022)—so sims can be used for sensitivity analysis based on many different counterfactuals. This can bolster generalizability and help researchers decide which sensitivity analyses to run with more costly human subjects studies.
As progress is made on these challenges, sims can be used in end-to-end complete studies when human subjects research is possible but impractical, such as with the limited funds of students and researchers in low-income countries, or fully impossible, such as a test of large-scale government policy change. Complete studies with sims can be partially validated through tractable human studies, such as testing individual components in a large-scale social system.
Each of these applications has a proximate goal, accurate findings that generalize current knowledge, and the five challenges span the potential inaccuracies as described in Table 1. Importantly, the use of sims does not circumvent many social science challenges; sims research still needs to be verifiable—we, as humans, have confidence in the results—and ethical, based on established standards such as the Belmont report (NCPHSBR, 1979).
@@ 6 Alternative Views
Because this is a new research area, there has been limited discussion of how promising sims are, but many of the reviewed works emphasize significant challenges and the limits of initial results (e.g., Gao et al., 2024; Wang et al., 2024a, b). Agnew et al. (2024) reviewed LLM simulation proposals, concluding that they “ignore and ultimately conflict with foundational values of work with human participants” due to problems such as diversity.
LLMs have been called “stochastic parrots” (Bender et al., 2021) that are “fundamentally not like us” (Shanahan, 2022) with “ineradicable defects” (Chomsky et al., 2023), among other critiques (Bender & Koller, 2020; LeCun, 2023; Marcus, 2022). These views imply that LLM capabilities are deeply limited, particularly on social simulation—a task that inherently requires the system to be humanlike or to understand humanlikeness sufficiently to make accurate predictions. We believe that theoretical arguments on this topic can only tell us so much, and we provide more detail on the limitations of our work in Appendix C. Ultimately, we believe that it will only be possible to validate the promise of LLM social simulations through rigorous empirical testing.
@@ 7 Conclusion
By developing an agenda for LLM social simulations (“sims”) with five tractable challenges, we show that the difficulties are not as insurmountable as some researchers have suggested. There is inevitably much uncertainty in the future of AI progress, but the rapid pace of technical advances requires us to take a step back and consider whether this endeavor could succeed. We argue that, by thoughtfully incorporating contemporary methods and future AI capabilities, sims are a promising research method to accelerate social science, open up new areas of inquiry into human behavior, and generate humanlike synthetic data to support the development of safe and beneficial AI.
Acknowledgments
[…]
We are grateful for feedback from members of the Stanford Human-Computer Interaction (HCI), Stanford Institute for Human-Centered AI (HAI), Stanford Digital Economy Lab (DEL), and University of Chicago Knowledge Lab. We particularly thank Suhaib Abdurahman, José Ramón Enríquez, Sophia Kazinnik, David Nguyen, and Joon Sung Park for their contributions.
Impact Statement
[…]
There are many potential societal impacts of our work. LLMs that effectively simulate human populations have the potential to advance scientific research, train human-compatible AI systems, and address a variety of other societal issues. However, they could be used irresponsibly, such as to generate online propaganda or train manipulative bots.
@@ A.1.2 Fundamental Limitations
There are fundamental problems that seem intractable even with an influx of funding. If researchers hope to collect real-world data, many of the most interesting research questions, such as the psychology of world leaders, the effects of large-scale policy change, or the effects of large-scale events on the general public (e.g., during a bank run; Kazinnik, 2023), would be logistically infeasible with any realistic amount of resources. Many social science questions are not even about real-world data, including future possibilities or past counterfactuals of the form “What would have happened if…?” Natural science has struggled less because physical laws and homogeneity allow for reliable laboratory testing. Human behavior has no chemical formula that can be isolated and studied in a lab.
@@ A.2 The Promise of LLM Social Simulations
This humanlikeness suggests new paradigms of where AI could be applied. Over seven decades since Alan Turing famously asked “Can machines think?” and proposed the Turing test (Turing, 1950), researchers at frontier AI labs now claim that current AI “attains a form of general intelligence” (Bubeck et al., 2023) and has reached “Level 1 General AI (‘Emerging AGI’)” (Morris et al., 2024). Generating realistic human behavioral data is the quintessential task that all humans can do, for at least one human—themselves. Thus, it would be reasonable to expect that in the next phase of AI systems, they would be able to do the same. This is especially the case as LLM pretraining is based on next-token prediction using trillions of words of human-written text, it is natural to expect high performance in simulating human behavior, at least that of reading and writing. As recent work has shown, LLMs have potential across a wide variety of domains, including economics (Horton, 2023), human-computer interaction (Hämäläinen et al., 2023), marketing (Brand et al., 2024), sociology (Kozlowski et al., 2024), political science (Argyle et al., 2023), and psychology (Abdurahman et al., 2024).
@@ A.2.3 Interview-Based Prompt Variation
Park et al. (2024a) conducted two-hour interviews with 1,052 participants, prepending them one at a time to the study information to create prompts for GPT-4o that would simulate each participant. To evaluate this method, they conducted an extensive battery of popular social science studies with each participant twice, two weeks apart. The difference between the human participant’s responses over the two weeks was used as the ceiling for simulation accuracy. Most strikingly, they found that the simulations were 85% as accurate in predicting responses to the General Social Survey [^2] https://gss.norc.org/ as participants’ first test results were at predicting their second test results, though the appropriate baseline is unclear because general knowledge—knowing the most common answers to typical survey questions, such as demographics—can perform much higher than random chance.
@@ Appendix C Limitations of Our Work
The literature on LLM simulations has only recently emerged and is scattered across disciplines and venues with varied terms and framings, so we were unable to identify work on the topic through conventional literature review methods (e.g., querying Scopus or Web of Science databases). We expect that we identified a majority of preprints and publications on the topic, given the consistency with which the authors have tended to become aware of new work, but this remains a concern—particularly for work that would be bibliometrically cut off from the aforementioned papers, such as if those unknown papers have not cited these works, if they are in different scholarly venues, or if they appear in non-English languages.
It is also difficult to draw a clear line between this topic and related topics, such as using LLMs for text annotation. For example, Veselovsky et al. (2023a) explicitly focus on the distributional accuracy of LLM-generated data relative to human data. The human data were from Abu Farha et al. (2022), social media posts made by crowdworkers that they self-identified as “sarcastic.” While Veselovsky et al. (2023a) have clearly compared LLM-generated and human-generated data, the data itself is more used for natural language processing research than for social science research with human subjects, so it has clear but limited relevance to our work. Likewise, our focus has been the simulation of humans in isolation, such as a survey context, but LLMs can also be used for multi-agent and multi-turn contexts (e.g., Park et al., 2023; Piao et al., 2025; Zhou et al., 2023). We maintain this focus in part because the success of more complex simulations will likely rely on the ability to simulate each human response in isolation.